OpenStudy (anonymous):

Ethan says there are a total of 700 hundreds in the number 7,000,000. Is Ethan correct? A. Yes, Ethan is correct. B. No, Ethan is not correct. There are exactly 700 ten thousands in 7,000,000. C. No, Ethan is not correct. There are exactly 7,000 hundreds in 7,000,000. D. No, Ethan is not correct. There are exactly 700 thousands in 7,000,000.

OpenStudy (anonymous):

B

OpenStudy (anonymous):

Yes, B is correct

jimthompson5910 (jim_thompson5910):

`700 hundreds = 700*100 = 70,000`

OpenStudy (thecatman):

i also say b 700 times 10000= 7000000

jimthompson5910 (jim_thompson5910):

`700 ten thousands = 700*10,000 = 7,000,000` `700 thousands = 700*1,000 = 700,000`